OpenStudy (anonymous):

Write in vertex form y=x²+3x-10 pleasee?

OpenStudy (saifoo.khan):

how many u have? a million?

OpenStudy (anonymous):

/: 10

OpenStudy (saifoo.khan):

\[(x+\frac32)^2-\frac{49}{4}\]

OpenStudy (saifoo.khan):

lol, i was kidding.

OpenStudy (saifoo.khan):

Thanks for fanning.

OpenStudy (anonymous):

lol okay and welcome (:

OpenStudy (saifoo.khan):

hehe.

OpenStudy (anonymous):

-3/2 = -1.5=h k=-12.25 vertex form= y= (x+1.5) - 49/4